# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

### v1

@@ III-C Average AoII under Linear Hierarchy
$\pi_{0_{c}}=\frac{p_{c}}{a},$ $\pi_{1_{m}}$ $=\frac{pp_{c}}{a(p+\rho^{-1})},$ (16)
$\pi_{0_{e}}=\frac{pp_{e}}{a},$ $\pi_{1_{u}}$ $=\frac{p^{2}\rho+pp_{e}}{a(p+\rho^{-1})},$
@@
For linear hierarchy scheme (5), the average AoII $x_{a}$ of a preemptive M/M/1/1 queue over the noisy channel at the monitor is
$x_{a}=\frac{m_{1}pp_{e}(1+\rho)}{a\lambda p_{c}}+\frac{(p_{e}+\rho)\big{[}pp_{c}+m_{2}(p^{2}\rho+pp_{e})\big{]}}{a\lambda p_{c}(p+\rho^{-1})},$ (17)
where $m_{1}\geqslant 1$ and $m_{2}\geqslant 1$ are slopes representing two different levels of dissatisfaction for freshness at state $0_{e}$ and $1_{u}$, respectively. Furthermore, if we set $m_{1}=m_{2}=1$, then (17) reduces to the simplest expression with only two levels,
@@
The derivation of (17) appears in Appendix B. The ergodicity of the SHSs Markov chain is established by (16). Hence, it is easy to verify that (17) always exists for any finite $m_{1}$ and $m_{2}$ based on the invertibility of matrix associated with (43), namely, $\det(\mathbf{B})=1-p_{e}>0$ for matrix $\mathbf{B}$ defined in (43). [^4] A general proof can be found in [9, Section V]. The simplified version (18) would later be used to compare with the AoI results in Section V.
Moreover, we note that (17) and (18) can be further specified to the results under metric AoS proposed in [21] by setting $p=1$. To be clearer, since the AoS metric is defined as the time since the content between two ends are out-of-sync, it is exactly the case of the proposed AoII metric where every newly sampled update changes its content with probability one. Hence, the expressions of average AoS in a preemptive M/M/1/1 queue over the noisy channel are given by
@@
III-D The Quadratic Hierarchy and Stability Conditions
The average AoII under the quadratic hierarchy is more complex since it involves a hybrid of two kinds of differential equations. The closed-form expression is summarized below.
@@ III-D The Quadratic Hierarchy and Stability Conditions
For quadratic hierarchy scheme (6), the average AoII $x_{b}$ of a preemptive M/M/1/1 queue over the noisy channel at the monitor is
@@
under certain stability conditions.
The derivation of (20) appears in Appendix C. However, we are more interested in the stability conditions of (20) since the invertibility of matrix $\mathbf{C}$ defined in (46) depends on the values of $k_{1}$ and $k_{2}$. These conditions are given below, the proof of which can be found in Appendix D.
@@
ii). The growth rates $k_{1}$ and $k_{2}$ at the quadratic growth states is limited, i.e., $0<k_{1}<\lambda<\mu$ and $0<k_{1}<k_{2}<\mu$.
Theorem 2 can be extended to other multi-level (more than four) hierarchies with finer state definitions by going through the same proof procedure as long as the hybrid only consists of linear and quadratic functions. For example, one can split the state $0_{e}$ into two states $0^{1}_{e}$ and $0^{2}_{e}$ representing different levels of error, where $0^{1}_{e}$ is defined the same as the original error state directly connected to $1_{m}$ and $0^{2}_{e}$ is defined as a severer error state that can only be reached by $1_{u}$. Meanwhile, we must point out that the above conditions are actually sufficient and necessary, as proved in Appendix D.
@@ IV SHSs Analysis over Collision Channel
In this section, we focus on another kind of unreliable channel that occurs when multiple nodes attempt to use the channel simultaneously. The SHSs analysis of the collision channel case is basically the same as the noisy channel case, except that the system state space $\mathbb{Q}_{c}$ is larger leading to a much more complex expression of the average AoII. Therefore, we only adopt the quadratic hierarchy scheme in this collision scenario considering the computation similarity and complexity. The stability conditions are also rederived.
IV-A SHSs Formulations with Quadratic Hierarchical AoII
We need to redefine the system states based on the collision nature of channel, where the state space $\mathbb{Q}_{c}$ consists of eight different states $\mathbf{q}\in\mathbb{Q}_{c}=\{0_{c},0_{e},1_{c},1_{m},1_{u},1_{e},2_{m},2_{u}\}$. States $0_{c},0_{e},1_{m}$ and $1_{u}$ are defined the same as the noisy channel case. However, we have to added four more states due to the existence of $X^{c}_{t}$, namely, $1_{c},1_{e},2_{m},2_{u}$. State $1_{c}$ is defined as the other correct (content-matched) state while the contender utilizes the channel to transmit $X^{c}_{t}$ alone. Hence, the AoII in $1_{c}$ would also not increase as in $0_{c}$. On the contrary, state $1_{e}$ is defined as the other error state while the contender utilizes the channel yet the content between ends is mismatched, where we assume that the AoII in $1_{e}$ grows the same as in $0_{e}$. States $2_{m}$ and $2_{u}$ are defined as two distinct collision states under different ageing processes when $X_{t}$ and $X^{c}_{t}$ both occupy the channel. To be clear, the former collision state $2_{m}$ can be directly reached by $1_{m}$ and $1_{c}$, which means the content of $\hat{X}_{t}$ at the monitor only lags one version behind the collision update $X_{t}$. However, the collision update $X_{t}$ in $2_{u}$ is considered to be more urgent for the same two reasons i) and ii) given in III-A, so the ageing process of $\hat{X}_{t}$ is severer than $2_{m}$ at the monitor.
@@ IV-A SHSs Formulations with Quadratic Hierarchical AoII
Based on the above state definitions, the quadratic hierarchy scheme considered here also has four levels, which is characterized by the state-dependent derivative $f_{3}^{\prime}(\mathbf{q},\mathbf{x},t)$.
- Quadratic hierarchy:
$f_{3}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},1_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},2_{m},\\ k_{1}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=0_{e},1_{e},1_{u},\\ k_{2}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=2_{u}.\\ \end{cases}$ (21)
The parameters $k_{1}$ and $k_{2}$ are the growth constants controlling the stability of SHSs, where $k_{1}<k_{2}$. Given that the collision channel is noise-free, state $2_{u}$ is assumed to be the only urgent state with the highest age dissatisfaction (the fourth level) since it is the only state that suffers from both collision and a severer out-of-sync situation. Note that this assumption can be easily generalized by adapting the growth rate constant $k_{1}$ and $k_{2}$.
@@
As we emphasized in the noisy channel case, the analysis and conclusion in this section could be extended to other multi-level hierarchies with finer state definitions. For example, one can track different urgent states in which successive arrivals of content-changed updates can lead to different levels of age dissatisfaction. However, these extensions are not trivial even in the two node case and each hierarchy scheme needs to be carefully designed due to the computation complexity of the SHSs approach, which can be indicated by our case below.
$\mathbf{q}=1_{u}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda$$\lambda_{c}$$\mu$$\mathbf{q}=1_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$q\lambda$$\mu$$\lambda_{c}$$p\lambda$$\mathbf{q}=1_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda_{c}$$\lambda$$\mu_{c}$$\mathbf{q}=1_{c}$$\mathbf{\dot{x}}=\mathbf{0_{n}}$$\lambda_{c}+q\lambda$$\mu_{c}$$p\lambda$$\mathbf{q}=2_{u}$$\mathbf{\dot{x}}=k_{2}\mathbf{x}$$\lambda_{c}+\lambda$$\mu_{c}$$\mu$$\mathbf{q}=2_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$\lambda_{c}+q\lambda$$\mu_{c}$$p\lambda$$\mu$$\mathbf{q}=0_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda$$\lambda_{c}$$\mathbf{q}=0_{c}$$\mathbf{\dot{x}}=\mathbf{0_{n}}$$q\lambda$$p\lambda$$\lambda_{c}$
@@
The SHSs states and corresponding transition set $\mathcal{L}_{c}$ are depicted in Fig. 4. For clarity, we now explain each transition $l\in\mathcal{L}_{c}$.
$l=0,6,9,13,16,20,24$: First, the assumption of the self-transition at state $0_{c}$ remains the same as the noisy channel case. In the remaining self-transitions, states $1_{c}$, $1_{m}$ and $2_{m}$ would stay unchanged when the transmission of updates is preempted either by a content-unchanged update (maintained) $X_{t}$ or the contender $X^{c}_{t}$. Moreover, states $1_{u}$, $1_{e}$ and $2_{u}$ would stay unchanged under any preemption.
@@
$1$ $0_{c}\to 1_{c}$ $\lambda_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$2$ $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
$3$ $0_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{0_{e}}&v^{2}_{0_{e}}\end{bmatrix}$
$4$ $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
$5$ $1_{c}\to 0_{c}$ $\mu_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$6$ $1_{c}\to 1_{c}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$7$ $1_{c}\to 2_{m}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{c}}&v^{2}_{1_{c}}\end{bmatrix}$
$8$ $1_{m}\to 0_{c}$ $\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$9$ $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
@@
$15$ $1_{e}\to 0_{e}$ $\mu_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$16$ $1_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$17$ $1_{e}\to 2_{u}$ $\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
@@
$19$ $2_{m}\to 1_{e}$ $\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
$20$ $2_{m}\to 2_{m}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
$21$ $2_{m}\to 2_{u}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
@@
$l=1,2$: Since a newly sampled update $X^{c}_{t}$ has no effect on the AoII of $X_{t}$ if the contender does not cause collision, the transition $l=1$ keeps $x_{1}=x_{2}=0$. However, A newly sampled update $X_{t}$ with changed content leads to mismatch between ends, and the AoII starts to grow subject to the linear (second level) age dissatisfaction after the state is transferred into $1_{m}$ by $l=2$.
$l=5,8,12$: If there is no collision, the content between ends would be correctly matched and the AoII is set to zero after a successful transmission corresponding to transitions $5,8,12$, respectively.
$l=11,21$: The update $X_{t}$ being transmitted is preempted by a new content-changed update. Hence, the mismatch between ends becomes severer and the states are transferred into $1_{u}$ and $2_{u}$ respectively. Since the preemption has no effect on the channel collision, the transition $l=21$ would keep the continuous state unchanged, i.e., $\mathbf{A}_{l}=\mathbf{I}$ is the identity matrix.
$l=7,10,14,17$: For transition $l=7$, the channel would collide if a newly sampled update $X_{t}$ with changed content arrives when the channel is currently busy at transmitting the update $X^{c}_{t}$ from contender. Similarly, for transitions $l=17$, the channel collides for any newly arrived update $X_{t}$ if $X^{c}_{t}$ is being transmitted. For transitions $l=10,14$, the channel would collide if any newly sampled update of the contender arrives when the channel is currently busy at transmitting the update $X_{t}$. The difference between $l=7,10$ and $l=14,17$ is in that the latter transfer the state into $2_{u}$, which suffers from the fourth level age dissatisfaction due to a severer out-of-sync situation between ends. Since the collision guarantees that neither update in transmission is successfully received, these four transitions also set $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$.
$l=18,19,22,23$: The channel becomes available from collision if either $X_{t}$ or $X^{c}_{t}$ is unsuccessfully delivered and discarded, where the transmission time of the left update is still exponential. The difference between $l=18,22$ and $l=19,23$ is in that the latter transfer the state into $1_{e}$, where no update of $X_{t}$ is being transmitted yet the mismatch of content still exists. Hence, $l=19,23$ lead to $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ instead of $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}0&x_{2}\end{bmatrix}$ in $l=18,22$.
$l=3,15$: The transitions $l=3,15$ between states $0_{e}$ and $1_{e}$ only depend on the contender source, which cannot lead to any possible drops on the AoII of $X_{t}$. Moreover, $x_{2}(t)$ keeps growing at the third level in these two error states because the content mismatch still exists. Hence, the continuous states are unchanged, i.e., $\mathbf{A}_{l}=\mathbf{I}$.
$l=4$: As assumed in the noisy channel case, any newly sampled update $X_{t}$ starts a transmission in $1_{u}$ when the system was in $0_{e}$ before.
@@
IV-B Average AoII under Quadratic Hierarchy
Following (12), the stationary distributions of the SHSs Markov chain in Fig. 4 are calculated as (LABEL:collision_distribution), where $m=\lambda_{c}+\lambda+\mu_{c}+\mu$ and $b=p(m\lambda+\lambda_{c}\mu)(m-q\lambda)+m\mu\mu_{c}$. [^5] The explicit expression of $\pi_{2_{u}}$ cannot be easily factorized, which is too long to be included in (LABEL:collision_distribution). Hence, we choose to present the implicit form.
Moreover, we could further obtain the average AoII of the collision channel case under a four-level hierarchy using Theorem 1.

### v2

@@ III-C Average AoII under Hierarchy Scheme 1
$\pi_{0_{c}}=\frac{p_{c}}{a},\$$\pi_{1_{m}}$ $=\frac{pp_{c}}{a(p+\rho^{-1})},$ (18)
$\pi_{0_{e}}=\frac{pp_{e}}{a},\$$\pi_{1_{u}}$ $=\frac{p^{2}\rho+pp_{e}}{a(p+\rho^{-1})},$
@@
For hierarchy scheme 1 in (7), the average AoII $x_{a}$ of a preemptive M/M/1/1 queue over the noisy channel is
$x_{a}=\frac{m_{1}pp_{e}(1+\rho)}{a\lambda p_{c}}+\frac{(p_{e}+\rho)\big[pp_{c}+m_{2}(p^{2}\rho+pp_{e})\big]}{a\lambda p_{c}(p+\rho^{-1})}.$ (19)
Furthermore, if we set $m_{1}=m_{2}=1$, then (19) reduces to the simplest expression with only two levels of hierarchy,
@@
The derivation of (19) appears in Appendix C. The ergodicity of the SHSs Markov chain is established by (18) and Lemma 2. Hence, it is easy to verify that (19) always exists for any finite $m_{1}$ and $m_{2}$ based on the invertibility of matrix $\mathbf{B}$ defined in (54). [^6] A general proof can be found in [6, Section V]. The simplified version (20) would later be used to compare with the AoI results in Section V.
Moreover, we note that (19) and (20) can be further specified to the results under metric AoS proposed in [18] by setting $p=1$. The AoS metric is defined as the time since the content between two ends are out-of-sync. This is exactly the case of the proposed AoII metric where every newly sampled update changes its content with probability one. Hence, the expressions of average AoS in a preemptive M/M/1/1 queue over the noisy channel are given by
@@
III-D Average AoII under Hierarchy Scheme 2
The average AoII under scheme 2 is more complex since it involves two kinds of ageing patterns. By leveraging Theorem 1, the closed-form expression is summarized below.
@@ III-D Average AoII under Hierarchy Scheme 2
For hierarchy scheme 2 in (8), the average AoII $x_{b}$ of a preemptive M/M/1/1 queue over the noisy channel is
@@
given that the steady-state stability conditions are satisfied.
The derivation of (22) appears in Appendix D. However, we are more interested in the steady-state stability conditions of (22), where the invertibility of matrix $\mathbf{C}$ defined in (57) depends on the values of $k_{1}$ and $k_{2}$. These conditions are given below, the proof of which can be found in Appendix E.
@@
ii). The growth constants $k_{1}$ and $k_{2}$ are limited, i.e., $0<k_{1}<\lambda<\mu$ and $0<k_{1}<k_{2}<\mu$.
Theorem 2 can be extended to other multi-level hierarchy scheme with finer state definitions as long as the hierarchy only consists of constant and linear growth rate functions. For example, one can split the state $0_{e}$ into two states $0^{1}_{e}$ and $0^{2}_{e}$ representing different levels of error, where $0^{1}_{e}$ is defined the same as the original error state directly connected to $1_{m}$ and $0^{2}_{e}$ is defined as a more severe error state that can only be reached by $1_{u}$. We also note that the above conditions are actually sufficient and necessary, as proved in Appendix E.
@@ IV SHSs Analysis over Collision Channel
In this section, we focus on another kind of unreliable channel that collides when multiple nodes attempt to use the channel simultaneously, in other words, the transmission time of these updates overlaps. The SHSs analysis is basically the same as the noisy channel case, except that the system state space $\mathbb{Q}_{c}$ is larger resulting in a much more complex expression of the average AoII. Therefore, under the collision channel, we only consider the hierarchy scheme with constant and linear growth rate functions since it is more representative.
In the following, we consider that only one contender source attempts to access the channel against the source of interest. The collision period ends if one of the updates finishes transmission first, which cannot be decoded and is discarded. Given that, there exists two different cases handling the remaining update on transmission: 1). It can be successfully decoded as long as there is no collision; 2). It cannot be successfully decoded. There is no fundamental difference between the analysis of case 1) and 2) by SHSs, except that the state spaces and transitions for two cases are slightly different. Therefore, we first analyze case 1) in the main body of this paper, while the discussion of case 2) is provided in the supplementary material due to the limitation of space.
Remark 4.
In practice, the model of collision 1) is suitable for the situation that the lengths, i.e., the transmission time, of two collided updates are significantly different. The collision period (overlap) is relatively short compared to the transmission time of the update with large transmission time. As a result, the effect of collisions on the content of remaining update is considered to be as small as possible. So the monitor is able to decode such updates under some advanced channel coding/decoding techniques. On the contrary, the model of collision 2) is suitable for the situation that the transmission time of two collided updates is close, which follows the collision model considered in [44]. Furthermore, since we allow preemption of updates for both sources, we assume that cases 1) and 2) are valid under preemption to facilitate the definition of discrete states in AoII SHSs. This assumption is only for the purpose of reducing the computation complexity of SHSs and demonstrating this work more clearly.
IV-A SHSs Formulations
For collision case 1), we first redefine the system states where the state space $\mathbb{Q}_{c}$ consists of eight different states $\mathbf{q}\in\mathbb{Q}_{c}=\{0_{c},0_{e},1_{c},1_{m},1_{u},1_{e},2_{m},2_{u}\}$. States $0_{c},0_{e},1_{m}$ and $1_{u}$ are defined the same as the noisy channel case. However, we have to added four more states due to the existence of $X^{c}_{t}$, namely, $1_{c},1_{e},2_{m},2_{u}$. State $1_{c}$ is defined as the other correct (content-matched) state while the contender utilizes the channel to transmit $X^{c}_{t}$ alone. Hence, the AoII in $1_{c}$ would also not increase as in $0_{c}$. On the contrary, state $1_{e}$ is defined as the other error state while the contender utilizes the channel yet the content between ends is mismatched, where we assume that the AoII in $1_{e}$ grows the same as in $0_{e}$. States $2_{m}$ and $2_{u}$ are defined as two distinct collision states with different AoII processes when the transmission of $X_{t}$ and $X^{c}_{t}$ overlaps. State $2_{m}$ can be directly reached by $1_{m}$ and $1_{c}$, which means the content of $\hat{X}_{t}$ at the monitor only lags one version behind that of $X_{t}$. However, the content of $X_{t}$ in $2_{u}$ is considered to be more urgent for the same two reasons i) and ii) given in III-A, so the AoII process of $\hat{X}_{t}$ is more severe than $2_{m}$.
@@ IV-A SHSs Formulations
Based on the above state definitions, the hierarchy scheme considered here also has four levels, which is characterized by the state-dependent growth rate function $f_{3}^{\prime}(\mathbf{q},\mathbf{x},t)$.
- Hierarchy Scheme 3:
[…]
With a slight notation abuse, the parameters $k_{1}$ and $k_{2}$ are also denoted as the growth constants controlling the steady-state stability of SHSs, where $k_{2}\geqslant k_{1}$. Given that the collision channel is noise-free, state $2_{u}$ is assumed to be the only urgent state with the highest age dissatisfaction (the fourth level) since it is the only state that suffers collisions and also a severe content-mismatch. Note that this assumption can be easily generalized by adapting the growth rate constant $k_{1}$ and $k_{2}$.
@@
As we emphasized in the noisy channel case, the analysis in this section can be extended to other multi-level hierarchy schemes using finer state definitions. For example, one can track different urgent states in which successive arrivals of content-changed updates can lead to different levels of age dissatisfaction. These extensions are not trivial even in the two node case. To that end, each hierarchy scheme needs to be carefully designed since the increasing of the number $M$ using finer state definitions makes the inverse of matrix in the AoII SHSs (very) hard to obtain, as shown in the case below.
[…]
@@
Positive Recurrence
The properties of positive recurrence and ergodicity in the collision channel case can be established by the same argument in the noisy channel case, hence, the proof here is omitted for simplicity.
The SHSs states and corresponding transition set $\mathcal{L}_{c}$ are depicted in Fig. 4. We now explain each transition $l\in\mathcal{L}_{c}$.
$l=0,6,9,13,16,20,24$: First, the assumption of the self-transition at state $0_{c}$ remains the same as the noisy channel case. In the remaining self-transitions, states $1_{c}$, $1_{m}$ and $2_{m}$ will stay unchanged when the transmission of updates is preempted either by a content-unchanged update $X_{t}$ or the contender $X^{c}_{t}$. Moreover, states $1_{u}$, $1_{e}$ and $2_{u}$ will stay unchanged under any preemption.
@@
11 $0_{c}\to 1_{c}$ $\lambda_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
22 $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
33 $0_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{0_{e}}&v^{2}_{0_{e}}\end{bmatrix}$
44 $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
55 $1_{c}\to 0_{c}$ $\mu_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
66 $1_{c}\to 1_{c}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
77 $1_{c}\to 2_{m}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{c}}&v^{2}_{1_{c}}\end{bmatrix}$
88 $1_{m}\to 0_{c}$ $\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
99 $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
@@
$15$ $1_{e}\to 0_{e}$ $\mu_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$16$ $1_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$17$ $1_{e}\to 2_{u}$ $\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
@@
1919 $2_{m}\to 1_{e}$ $\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
2020 $2_{m}\to 2_{m}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
2121 $2_{m}\to 2_{u}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
@@
$l=1,2$: Since a newly-sampled update $X^{c}_{t}$ has no effect on the AoII of $X_{t}$ without collisions, the transition $l=1$ keeps $x_{1}=x_{2}=0$. However, A newly-sampled update $X_{t}$ with changed content leads to mismatch, and the AoII starts to grow after the state is transferred into $1_{m}$ by $l=2$.
$l=5,8,12$: The AoII is reset to or remain zero after a successful transmission.
$l=11,21$: The update $X_{t}$ being transmitted is preempted by a new content-changed update. Hence, the mismatch becomes severe. The states are transferred into $1_{u}$ and $2_{u}$ respectively. Note that the preemption has no effect on the channel collision in the transition $l=21$.
$l=7,10,14,17$: When the channel is currently busy at transmitting the update $X^{c}_{t}$, collisions arise if a newly-sampled update $X_{t}$ with changed content arrives, i.e., $l=7$. Similarly, for $l=17$, the channel collides for any newly-arrived update $X_{t}$ when $X^{c}_{t}$ is being transmitted. For $l=10,14$, collisions arise if any newly-sampled update of the contender arrives when the channel is currently busy at transmitting the update $X_{t}$. The difference between $l=7,10$ and $l=14,17$ is in that the latter transfers the state into $2_{u}$, which suffers the fourth level age dissatisfaction due to a more severe content mismatch. Since neither update in transmission can be successfully decoded during the collision period, these four transitions all set $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$.
$l=18,19,22,23$: The collision ends if either $X_{t}$ ($l=18,22$) or $X^{c}_{t}$ ($l=18,23$) finishes transmission. The difference between $l=18,22$ and $l=19,23$ is in that the latter transfers the state into $1_{e}$, where no update of $X_{t}$ is being transmitted yet the mismatch of content still exists. This leads to $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ instead of $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}0&x_{2}\end{bmatrix}$ in $l=18,22$.
$l=3,15$: The transitions $l=3,15$ between states $0_{e}$ and $1_{e}$ only depend on the contender source, which cannot change the AoII of $X_{t}$. Moreover, the AoII keeps growing at the third level in these two error states because the content mismatch still exists.
$l=4$: If the system were in the error state $0_{e}$, any newly-sampled update $X_{t}$ transfers the system state into $1_{u}$ and starts a transmission.
@@
IV-B Average AoII under Collision Channels
For collision case 1), the stationary distributions of the SHSs Markov chain in Fig. 4 are calculated as (LABEL:collision_distribution), where $m=\lambda_{c}+\lambda+\mu_{c}+\mu$ and $b=p(m\lambda+\lambda_{c}\mu)(m-q\lambda)+m\mu\mu_{c}$. [^7] The explicit expression of $\pi_{2_{u}}$ cannot be easily factorized, which is too long to be included in (LABEL:collision_distribution). Hence, we choose to present the implicit form.
Based on these, we obtain the average AoII of the collision channel case under a four-level hierarchy using Theorem 1.